OpenStudy (anonymous):

find the value of b in the equation 25^b = 5square root 5

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

actually I want to thank you for this question, It refreshed my memory.

OpenStudy (anonymous):

25^b = 5(sq.rt.5) log (25)^b = log 5(sq.rt.5) b*log (25) = log 5(sq.rt.5) b= (log 5(sq.rt.5)) / (log (25) b=.75